OpenStudy (anonymous):

f(x) = 3x + 1; g(x) = 5x - 1 Find f/g.

OpenStudy (anonymous):

I would like an explanation:)

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

i.e. write a big fraction bar, put \(f(x)\) on top and \(g(x)\) on the bottom

OpenStudy (anonymous):

So, 3x+1/5x-1? with what restriction?

OpenStudy (anonymous):

\[(f/g)(x) = \frac{f(x)}{g(x)}\]

OpenStudy (anonymous):

yes, like that restriction is that the denominator cannot be zero, so write \[5x-1=0\] and solve for \(x\)

OpenStudy (anonymous):

1/5....But how'd you get that equation?

OpenStudy (anonymous):

the denominator is \(5x-1\) right?

OpenStudy (anonymous):

oh! right. Thanks!:)

OpenStudy (anonymous):

since you are not allowed to divide by zero (division by zero is undefined) i set it equal to zero

OpenStudy (anonymous):

yw